A stone whose mass is 0.1 kg rests on a horizontal frictionless surface. A bullet of mass 0.0025 kg, traveling horizontally at 400 m/s, strikes the stone and rebounds horizontally at right angle to its original direction with a speed of 300 m/s. Compute the magnitude and direction of the velocity of the stone after it is struck


1
Expert's answer
2021-05-26T21:35:06-0400

Gives

"m_1=0.0025kg"

"m_2=0.1kg"

"u_1=400m\/sec"

"u_2=0m\/sec"

"v_1=300m\/sec"

X-axis components


"m_1u_1+m_2u_2=m_1u_1cos\theta+m_2u_2cos\phi"

"0.0025\times400+0.1\times0=0.0025\times300cos90°+0.1\times v_2cos\phi"

"1=0+0.1v_2cos\phi"

"v_2cos\phi=10\rightarrow(1)"

Y -axis components

"0+0=m_1v_1sin\theta-m_2v_2sin\phi"



"0=0.0025\times 300\times sin90°-0.1\times v_2sin\phi"

"0.75-0.1v_2sin\phi"

"v_2sin\phi=7.5\rightarrow(2)"

Eqution (2)/equation (1)

"\frac{v_2sin\phi}{v_2cos\phi}=\frac{7.5}{10}"

"tan\phi=0.75"

"\phi=36.86°"

Put value in eqution (1)

"10=v_2cos(36.86\u00b0)"

"v_2=12.49m\/sec"